OpenStudy (anonymous):

Solve for y. 9x-3y=12 ( literal equation)

OpenStudy (mathstudent55):

First subtract 9x from both sides. Then divide both sides by -3.

OpenStudy (anonymous):

I got y=-3x+4 but it is still wrong. Thank you btw

OpenStudy (anonymous):

9x-3y=12 -9x -9x -3y=-9x+12 (-3y)/-3=(-9x+12)/-3 y=3x-4

OpenStudy (anonymous):

You divided each side by 3, instead of -3.

OpenStudy (anonymous):

Oohh ok I understand now. thank you very much ! God bless

OpenStudy (anonymous):

No problemo! :)